Paul Martin Walker, the credit analyst at the Washington Mutual bank received a credit application from Seether, LLC, a private company. An abbreviated portion of the financial information provided by the company is shown below:

Total assets

EBIT

Net working capital

Book value of equity

Accumulated retained earnings

Sales

Total liabilities

$63,000

$7,900

$4,200

$18,000

$16,000

$92,000

$57,000

What is the Z-score for this company?

Select one:

A. 3.519

B. there is not enough information to calculate Seether’s Z-score

C. 2.851
Answer #1

Z-score is given by:

Z-score = 6.56*Net working capital/total assets + 3.26*Accumulated retained earnings/total assets + 1.05*EBIT/total assets + 6.72*Book value of equity/total liabilities

= (6.56*4,200/63,000) + (3.26*16,000/63,000) + (1.05*7,900/63,000) + (6.72*18,000/57,000) = 3.519 (Option A)